willienewton22:

An empty box is shaped like a rectangular prism. The box has a base area of `9/10` square foot and a height of `1/3` foot. How much packing material is required to fill the box?

jhonyy9:



jhonyy9:

total lateral area \[A_{lateral} = 4×\frac{ 1 }{ 4 } = \frac{ 4 }{ 4 }= 1 ft^2\] \[A_{total} = A_{lateral} + A_{base} + A_{topside}\] \[A_{total} = 1+\frac{ 9 }{ 10 } +\frac{ 9 }{ 10 }= \frac{ 28 }{ 10 } = \frac{ 14 }{ 5 }= 2\frac{ 4 }{ 5 }ft^2\]

jhonyy9:

@AZ opinion pls - ty